# Exact Chiral Symmetries of 3+1D Hamiltonian Lattice Fermions

### v1

@@ I Introduction
Regulating chiral gauge theories like the standard model on the lattice has been a long standing problem. As well as offering a route for extracting more precise predictions from these theories, this has been highlighted as a deep theoretical problem thanks to early no-go results [1, 2]. These prove under certain assumptions that these lattice theories always come with fermions of both handedness, a phenomenon known as fermion doubling. This problem already exists for fermions with chiral global symmetries, and can be understood as a ’t Hooft anomaly [3]. More precisely, a system with an on-site global symmetry, meaning one which does not mix degrees of freedom at separate sites, must be free of ’t Hooft anomalies. This rules out, for example, a lattice model with an on-site $U(1)$ global symmetry giving rise to a single charged Weyl fermion in the infrared (IR).
There have been several approaches so far attempted to circumvent the fermion doubling problem. A simple strategy is to try to gap out the unwanted fermions. Because of the above, any such perturbation also breaks the chiral symmetry. However, it was observed by Ginsparg and Wilson [4] that in Euclidean spacetime lattice models, there is a kind of remnant of the chiral symmetry. This is in fact another symmetry [5], which takes the form
@@
where $a$ is the lattice spacing, and $D$ is a finite difference operator which is a discretization of the spacetime Dirac operator. This is an example of a not-on-site symmetry, since $(D\psi)(x)$ is a (linear) function of $\psi$ at nearby spacetime lattice points, and not of $\psi(x)$ alone. When $a\to 0$ this becomes the continuum axial symmetry of a Dirac fermion. It thus protects the fermion from perturbatively gaining a mass.
In this paper, we will construct Hamiltonian models on a 3d cubic lattice with chiral global symmetry. These models are Weyl semimetals [6, 7] protected by not-on-site symmetries, and have the advantage over Euclidean lattice models of being realizable in quantum simulators [8, 9].
In particular, we will discuss a model which in the IR gives rise to a single Weyl fermion, and in the ultraviolet (UV) has a not-on-site symmetry which flows to the chiral $U(1)$ symmetry, and protects the Weyl fermion from obtaining a Majorana mass.
Then we will discuss a model containing two Weyl fermions, which is in fact a simple model of a magnetic Weyl semimetal, and show that it has not-on-site symmetries which flow to the chiral $SU(2)$ symmetry of the infrared. This protects both Weyls from obtaining a mass, thanks to Witten’s $SU(2)$ anomaly [10]. This insight offers a very different interpretation of the usual chiral anomaly involving $U(1)$ and translation [11, 12, 13], as such an anomaly holds even when crystalline translation symmetry is broken. We develop a formalism which allows one to explore general not-on-site symmetries acting linearly on the fermion operators, analogous to the Ginsparg-Wilson symmetry.
Finally, we will discuss time reversal symmetry in these systems, and build a 2+1d model with a single massless Dirac fermion in the IR, protected by the parity anomaly.
@@ II Symmetry-protected single Weyl fermion in 3+1d
$h_{2}^{\mathrm{BdG}}(\mathbf{k})=$ $\frac{1}{2}\left[\sin k_{x}\sigma^{x}+\sin k_{y}\sigma^{z}+\sin k_{z}\tau^{z}\sigma^{y}+m(\mathbf{k})\sigma^{y}\right],$ (4)
@@
We see that, like the Ginsparg-Wilson symmetry, this charge operator involves nearest neighbor couplings, and is thus not-on-site
@@ III Symmetry-protected double Weyl fermion in 3+1d
The first is the usual $U(1)$ symmetry, while the second is composed of Kitaev Majorana chains [16] along $z$-axis wires. These generators do not commute, instead they generate an infinite-dimensional Lie algebra known as the Onsager algebra, introduced in [17]. This algebra has recently appeared in the study of the 1+1D chiral anomaly on the lattice [18, 19], and offers another route to defining exact symmetries on the lattice giving anomalous symmetries in the IR.
@@ IV Time Reversal Symmetric Single Weyl Fermion in 3+1d
$B(\mathbf{k},w)=\begin{cases}e^{-\frac{w^{2}}{|\mathbf{k}|^{2}-w^{2}}}&\mathrm{for}\,\,|\mathbf{k}|<w\\ 0&\mathrm{for}\,\,|\mathbf{k}|\geq w\end{cases}$ (16)
@@ VI Discussion
The chiral symmetry in 3+1d has a nearest-neighbor charge operator, but is non-quantized, as it must be by the no-go theorem of [14] and our arguments above. The two $SU(2)$ generators we constructed are quantized (and one is on-site), but don’t satisfy the expected $SU(2)$ Lie algebra, instead satisfying an Onsager algebra. This is consistent with the anomaly since either $U(1)$ symmetry on its own is anomaly-free. In this way, it is structurally similar to the $U(1)_{V}\times U(1)_{A}$ anomaly realized in 1+1D in [18, 19], where both generators are quantized and one is on-site, but they don’t commute.
@@
Since Weyl semimetals are physically realizable in a variety of condensed matter experiments [24, 25] it may be interesting to explore potentially observable effects of this low-energy $SU(2)$ symmetry and associated anomaly in such systems, although the global nature of the anomaly may complicate things.
Acknowledgements: LG acknowledges support from the Walter Burke Institute for Theoretical Physics at Caltech and the Caltech Institute for Quantum Information and Matter. RT acknowledges support from the Mani L. Bhaumik Presidential Term Chair in Theoretical Physics at UCLA.

### v2

@@ I Introduction
Regulating chiral gauge theories like the standard model on the lattice has been a long standing problem. As well as offering a route for extracting more precise predictions from these theories, this has been highlighted as a deep theoretical problem thanks to early no-go results like the Nielsen-Ninomiya theorem [1, 2]. These prove that lattice theories with certain features always come with fermions of both handedness, a phenomenon known as fermion doubling. This problem already exists for fermions with chiral global symmetries, and can be understood as a consequence of ’t Hooft anomaly matching [3]. In particular, a system with an on-site global symmetry, meaning one which does not mix degrees of freedom at separate spacetime points, must be free of ’t Hooft anomalies in the infrared (IR). This rules out, for example, a 3+1D lattice model with an on-site $U(1)$ global symmetry giving rise to a single charged Weyl fermion at low energies.
There have been several approaches so far to circumvent the fermion doubling problem. A simple strategy is to try to “gap out” the unwanted fermions by introducing carefully chosen mass terms. Examples of this construction are Wilson fermions [4] and Kogut-Susskind fermions [5]. These fermions are commonly used in lattice simulations of QCD, whose gauge symmetry is vector-like (non-chiral) and can be realized in an on-site fashion, allowing coupling to gauge fields.
Because of anomaly matching, anomalous chiral flavor symmetries of these models are explicitly broken by the mass terms. This puts the fermions in danger of being fine tuned, unless there is enough remnant symmetry to protect them from obtaining other mass terms (i.e. to prevent additive mass renormalization). This usually comes in the form of discrete crystalline symmetries of the lattice model such as translation, which can act chirally. This mechanism also stabilizes gapless fermions in condensed matter systems such as Weyl semimetals [6, 7, 8, 9, 10].
Another method to stabilize such fermions is to use more general not-on-site symmetries, which can give rise to continuous chiral symmetries. An early example is the Ginsparg-Wilson symmetry of a Euclidean lattice fermion [11, 12] which infinitesimally takes takes the form
@@
where $a$ is the lattice spacing, and $D$ is a finite difference operator which is a discretization of the spacetime Dirac operator. The operator $(D\psi)(x)$ is a (linear) function of $\psi$ at nearby spacetime lattice points, which makes it not-on-site, whereas an on-site symmetry would be a function of $\psi(x)$ alone (and therefore anomaly-free). When $a\to 0$ this becomes the continuum axial symmetry of a Dirac fermion. It thus protects the fermion from gaining a mass. To emphasize that this axial symmetry is not just emergent, but that it is forced by the UV Ginsparg-Wilson symmetry, we call it an “emanant symmetry” after [13].
In this paper, we will construct Hamiltonian models with minimal number of fermions at low energy, transforming anomalously under not-on-site symmetries. In particular, we construct three main models
  1. 1.
    a 3+1D model with a single Weyl fermion having an emanant $U(1)$ chiral symmetry
  2. 2.
    a 3+1D model with a pair of Weyl fermions having an emanant $SU(2)$ chiral symmetry with Witten’s global $SU(2)$ anomaly [14]
  3. 3.
    a 2+1D model with a single Dirac fermion and an emanant $U(1)\rtimes T$ symmetry with the parity anomaly
We emphasize that these models have the minimal number of fermions for their emanant anomalous symmetries, and that each is protected from gaining a mass—they are not fine tuned. Other Hamiltonian families of models of fermions enjoying Ginsparg-Wilson symmetries exist, such as overlap fermions [15, 16, 17, 18]. As far as we know, our constructions give the first with these minimal representations and emanant continuous symmetry groups.
From the condensed matter point of view, these are models of Weyl semimetals [6, 7]. Using band theory, and especially the Bogoliubov-de-Gennes (BdG) formalism, we are able to study a large class of transformations generalizing the Ginsparg-Wilson symmetry (1) which are linear in the fermion operators, which we leverage to produce the models. We can also derive some no-go results in this formalism, constraining what these not-on-site symmetries must look like for certain anomalous symmetries to emanate from them, making contact with other no-go theorems generalizing Nielsen-Ninomiya such as in [19].
@@ II Symmetry-protected single Weyl fermion in 3+1d
$h_{2}^{\mathrm{BdG}}(\mathbf{k})=$ $\frac{1}{2}\left[\sin k_{x}\sigma^{x}+\sin k_{y}\sigma^{z}+\sin k_{z}\tau^{z}\sigma^{y}+m(\mathbf{k})\sigma^{y}\right].$ (4)
@@
We see that, like the Ginsparg-Wilson symmetry, this charge operator involves nearest neighbor couplings, and is thus not-on-site.
@@ III Symmetry-protected double Weyl fermion in 3+1d
The first is the usual $U(1)$ symmetry, while the second is composed of Kitaev Majorana chains [21] along $z$-axis wires. These generators do not commute, instead they generate an infinite-dimensional Lie algebra known as the Onsager algebra, introduced in [22]. This algebra has recently appeared in the study of the 1+1D chiral anomaly on the lattice [23, 24, 25], and offers another route to defining exact symmetries on the lattice giving anomalous symmetries in the IR.
@@ IV Time Reversal Symmetric Single Weyl Fermion in 3+1d
$B(\mathbf{k},w)=\begin{cases}e^{\frac{|\mathbf{k}|^{2}}{|\mathbf{k}|^{2}-w^{2}}}&\mathrm{for}\,\,|\mathbf{k}|<w\\ 0&\mathrm{for}\,\,|\mathbf{k}|\geq w\end{cases}$ (16)
@@ VI Discussion
The chiral symmetry in 3+1d has a nearest-neighbor charge operator, but is non-quantized, as it must be by the no-go theorem of [19] and our arguments above. The two $SU(2)$ generators we constructed are quantized (and one is on-site), but do not satisfy the expected $SU(2)$ Lie algebra, instead generating an infinite-dimensional Onsager algebra. This is consistent with the anomaly since either $U(1)$ symmetry on its own is anomaly-free. In this way, it is structurally similar to the $U(1)_{V}\times U(1)_{A}$ anomaly realized in 1+1D in [24, 25], where both generators are quantized and one is on-site, but they don’t commute.
@@
Finally, we would eventually like to make lattice models of chiral gauge theories. Although we have produced models with chiral symmetries, the not-on-siteness of the symmetries makes them difficult to gauge. In fact one expects on general grounds that because of the nonvanishing ’t Hooft anomalies, we should not be able to gauge these symmetries. For interesting anomaly cancellation, such as in the Standard Model, one needs low energy fermions of different charges. It is not clear models of the kind we studied can produce such charge assigments. It would be very interesting if in some more complicated models with not-on-site chiral symmetries having vanishing ’t Hooft anomalies, if we can nonetheless find some way to gauge them, as we done for discrete symmetries in 1+1D in [30].
Acknowledgements: We are grateful for conversations with David Kaplan, Shu-Heng Shao, and Cenke Xu. LG acknowledges support from the Walter Burke Institute for Theoretical Physics at Caltech and the Caltech Institute for Quantum Information and Matter. RT acknowledges support from the Mani L. Bhaumik Presidential Term Chair in Theoretical Physics at UCLA.